OpenStudy (anonymous):

Find an equation of the line passing through the point (5,4) and parallel to the line whose equation is 2x+y=3

OpenStudy (anonymous):

y=-2x+14

OpenStudy (anonymous):

How did you get that solution?

OpenStudy (anonymous):

y=-2x+3 keep the slope. substitute x and y values. 4=-2(5)=b this is to find the new equation 4=-10+b , b=14 replace b in the original y=-2x=14

OpenStudy (anonymous):

Thanks!